Xania Inc. uses a normal job-order costing system. Currently, a plantwide overhead rate based on machine hours is used. Xania's plant manager has heard that departmental overhead rates can offer significantly better cost assignments than a plantwide rate can offer. Xania has the following data for its two departments for the coming year:

Department A

Department B

Overhead costs (expected)

$750,000  

$330,000  

Normal activity (machine hours)

  6,000  

    10,000  

Required:

  1. Compute a predetermined overhead rate for the plant as a whole based on machine hours.
  1. Compute predetermined overhead rates for each department using machine hours. (Note: Carry your calculations out to the nearest dollar.)

  1. Job 73 used 25 machine hours from Department A and 50 machine hours from Department B. Job 74 used 40 machine hours from Department A and 15 machine hours from Department B. Compute the overhead cost assigned to each job using the plantwide rate computed in Requirement 1. Repeat the computation using the departmental rates found in Requirement 2.

Job 73 Plantwide

Job 73 Departmental

Job 74 Plantwide

Job 74 Departmental

  1. Using a Plantwide, if actual machine hours for the year were 17,000 hours and actual overhead costs were $1,100,000, how much was overhead under- or over-applied for the year and how would that affect Cost of Goods Sold when it is closed to that account.

$$_________________________ over-applied/under-applied (circle one)

$$_________________________ Cost of Goods Sold increased/decreased (circle one)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Xania Inc.
Answer 1
Calculation of Predetermined OH Rate for the plant Department A Department B Total Note
Machine Hours              6,000.00            10,000.00        16,000.00 A
Overhead costs (expected)          750,000.00         330,000.00 1,080,000.00 B
Predetermined OH Rate for the plant                67.50 C=B/A
Answer 2
Calculation of Predetermined OH Rate for each department Department A Department B
Machine Hours              6,000.00            10,000.00 D
Overhead costs (expected)          750,000.00         330,000.00 E
Predetermined OH Rate for each department                 125.00                   33.00 F=D/E
Answer 3
Overhead cost assigned to Job 73 Department A Department B Total
Machine Hours used                    25.00                   50.00                75.00 G
Predetermined OH Rate for the plant                67.50 See C
Overhead cost assigned to Job 73          5,063.00 H=G*C
Overhead cost assigned to Job 74 Department A Department B Total
Machine Hours used                    40.00                   15.00                55.00 I
Predetermined OH Rate for the plant                67.50 See C
Overhead cost assigned to Job 74          3,713.00 J=I*C
Answer 4
Overhead cost assigned to Job 73 Department A Department B Total
Machine Hours used                    25.00                   50.00 See G
Predetermined OH Rate for the department                 125.00                   33.00 See F
Overhead cost assigned to Job 73              3,125.00              1,650.00          4,775.00 K=G*F
Overhead cost assigned to Job 74 Department A Department B Total
Machine Hours used                    40.00                   15.00 See I
Predetermined OH Rate for the department                 125.00                   33.00 See F
Overhead cost assigned to Job 74              5,000.00                 495.00          5,495.00 L=I*F
Answer 5
Total Machine Hours            17,000.00 M
Predetermined OH Rate for the plant                    67.50 See C
Applied overhead      1,147,500.00 N=M*C
Actual overhead costs      1,100,000.00 O
Over applied Overhead            47,500.00 P=N-O
Journal Entry
Account Debut $ Credit $
Overhead costs (Over applied)            47,500.00 See P
Cost of Goods Sold            47,500.00 See P